OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)).

OpenStudy (aakashsudhakar):

You're given that f(x) = 7 - 4x and g(x) = 10x - 6. f(g(x)) = f(10x - 6) = 7 - 4(10x - 6) = 7 - 40x + 24 = 31 - 40x f(g(x)) = 31 - 40x

OpenStudy (shamil98):

f(g(x)) = ( f o g ) you input g(x) into the equation for f(x) So what would you do?

OpenStudy (anonymous):

-4(10x-6)+7= -40x+24+7=31-40x